a student received grades of 80, 85, and 76 on three exams. find the least grade that the student must get on the fourth exam to obtain an average of no less than 84.
Answer by rfer(12662)

(80+85+76+x)/4=84
241+x=84*4
241+x=336
x=336-241
x=95